OpenStudy (donnie1999):

From Y=1/2 lxl + 4/5 , what are the values of a, h, and k?

OpenStudy (kaisertheslayer):

what are the choices?

OpenStudy (donnie1999):

a. A=2, h=0, k=4/5 b. A=2, h=4, k=5 c. A=1/2, h=0, k=1 d. A=1/2, h=0, k=4/5 e. A=1, h=0, k=0

OpenStudy (donnie1999):

@KaiserTheSlayer

OpenStudy (caominhim):

what do A, H, and K represent?

OpenStudy (donnie1999):

idk

OpenStudy (caominhim):

that's kind of a problem

OpenStudy (caominhim):

nobody's going to be able to solve this until we know what it is the problem is asking

OpenStudy (kaisertheslayer):

it would be D.

OpenStudy (kaisertheslayer):

because a.=1/2 h.=x coordinate of vertex and k.= y coordinate of vertex

OpenStudy (caominhim):

how did you decide that's what the values were?

OpenStudy (kaisertheslayer):

because a. is the number infant of the variable

OpenStudy (kaisertheslayer):

and then h and k are the coordinates of the vertex

OpenStudy (caominhim):

and how did you come to that conclusion? is that a general fact at your school?

OpenStudy (caominhim):

if i was given 3 variables with no attribution I would have had to request clarification

OpenStudy (caominhim):

@KaiserTheSlayer and what is a number infant?

OpenStudy (donnie1999):

@caominhim can u stop commenting on my questions because ur not helping me.

OpenStudy (caominhim):

did you learn from this thread? can you now do a different problem of the same type and understand what you're doing?

OpenStudy (donnie1999):

can u leave me alone.